During a tornado in 2008 the peachtree plaza westin hotel in downtown atlanta suffered damage. Suppose a piece of glass dropped near the top of the hotel falling 215 meters? Ignoring air resistance, how long would it take the piece of galss to hit the ground? Ignoring air resistance, what will the velocity of the piece of glass be when it strikes the ground?
1
Expert's answer
2020-09-21T06:29:03-0400

The time it would take for the glass to reach the ground:


"t=\sqrt{\frac{2h}{g}}=6.62\text{ s}."

Ignoring air resistance, the velocity near the ground will be


"v=\sqrt{2gh}=64.9\text{ m/s}."